40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10372     Entry time: Wed Aug 13 03:03:37 2014
Author: ericq 
Type: Update 
Category: General 
Subject: Green beatnote troubles 

[Jenne, Rana, ericq]

No luck locking tonight, as spent a while trying to figure out the complete absence of the green beatnotes. Long story short, we ended up having to adjust the pointing on the PSL table.

Unrelated to this, we also turned on the noise eater on the PSL laser because why not. 


We hooked the BBPDs directly up to a 300MHz scope to try to see the beat as it happened. We witnessed a very strange intermittent ~800MHz oscillation on the Y BBPD, and weirder still, on both the RF and DC outputs of the PD, and the frequency was independent of the laser temperatures. This is to be investigated in the future, but was not related to the beat note state. 

Some progress was made when we took some components out, and looked at the far field of the PSL-Ygreen overlap, and saw some misalignment, and corrected it. Putting the end laser temperature in the usual area allowed the beat note to be found, with the eventual amplitude of ~-40dBm directly out of the BBPD. The Y green alignment was pretty bad throughout, so this can be improved to bring the beat amplitude up. We should also check and make sure we're well aligned to the SHG with the PSL light. We're leaving the X beat for tomorrow, now knowing that we should be able to get it with careful alignment. 

ELOG V3.1.3-